an iron rode 1 meter long is to be cut in to two parts so that 1/2 of one equal to 1/3 of the other
Answer by skartikey(21)

let two parts are x and 1-x
we are given
(1/2)x=(1/3)(1-x)
3x=2(1-x)
3x=2-2x
5x=2
x=0.4
1-x=1-0.4=0.6
two parts are 0.4 meter and 0.6 meter